# Intrinsic decay rates and steady states of driven Josephson junction chains cavities

### 06 Mar 2026 v1

@@ I Introduction
Figure 1: (a) Sketch of system, consisting of a chain of JJs coupled to a transmission line. Small squares denote adjacent superconducting islands. Lines in different shades of magenta color illustrate the wave function of the first plasmonic modes with number $k=1,2,3$. The coupling to terminals $\kappa_{\text{ex}}$ and to the internal environment $\kappa_{\text{i}}$ are also shown. (b) Detail of the circuit, with $E_{J}$ the Josephson energy, $C_{g}$ the capacitance to ground, and $\theta_{i}$ the superconducting phase of each island. (c) Comparing exact, Eq. (6), and expanded, Eq. (11) dispersion relation of the chain shows agreement for low-lying modes.
@@ II.1 From the non-interacting Hamiltonian to four-wave mixing
The first term is the charging energy and the second one is the nonlinear Josephson potential, with $E_{J}=(\hbar/2e)^{2}/L_{J}$, $L_{J}$ being the inductance of a single junction. The two variables $Q_{n}$ and $\theta_{n}$ are canonically conjugated, $\theta_{n},Q_{m}]=\mathrm{i}2e\delta_{n,m}$. Due to the typically large impedance mismatch between the chain and the surrounding network, we are assuming $\theta_{0}=\theta_{N}=0$. Notice that such constraint is responsible for the limits 1 and N2N-2 of the second sum in Eq. (1). Regarding the charging part, $Q_{n}$ are the canonical variables conjugate to the phases, but it is sufficient to define them only for the phases that are actual dynamical degrees of freedom, namely $\theta_{1},\ldots,\theta_{N-1}$, given the others are set to zero. For this reason, the charging sum runs from $1$ to N1N-1. In the simplest case, each junction is only coupled to ground with on-site capacitance $C_{g}$ and to its nearest neighbors with capacitance $C_{J}$, such that the capacitance matrix $C_{nm}$ is tridiagonal,
@@
In the remainder of the paper, we will neglect Kerr corrections and rather focus on four-wave mixing. We restricted to number-conserving terms as previous works proved them to play a dominant role in the relaxation of bosons in similar contexts [48, 49]. Also, differently from two-into-two scattering, the one-into-three scattering does not have on-shell configurations around which the energy violation is weak, resulting in unfavorable off-resonant conditions. More details can be found in Appendix C.
@@ II.2 Kinetic equation and decay rates
To describe the scattering processes from quartic nonlinearity, we use the framework of a kinetic equation [50, 51, 52]. The kinetic equation provides a simplified description of the interacting modes, that neglects coherences between different modes and assumes that the state of the system is specified by the occupation numbers of individual modes,
@@ II.3 Relevant parameters and cavity language
Parameter Scaling Typical value
$\mathcal{F}$ $\frac{\sqrt{E_{g}E_{J}}}{N\kappa^{0}}$ $\approx 14$
$\zeta$ $\frac{kE_{g}}{N^{2}E_{c}}$ $\sim 10^{-5}-10^{-3}$
$\bar{p}$ (w/o drive) $\frac{TN}{\sqrt{E_{g}E_{J}}}$ $\approx 30$
$g$ $\frac{1}{N^{2}}\sqrt{\frac{E_{g}}{E_{J}}}$ $\approx 1.8\cdot 10^{-8}$
Table 2: When appearing in the above estimates, we use $T=0.1\,\text{K}$ and k=2200k=2-200.
The parameters discussed before can be thought as intrinsic properties of JJ chains. Now, we discuss two parameters that are strongly influenced by the presence of driving. The first parameter is (effective) temperature in units of $\Delta_{1}$, that determines how many modes are occupied by at least one photon,
@@ III.1 Different processes contributing to decays
where an expansion assuming $\xi\ll 1$ was performed. For realistic values of parameter $\xi$, that are typically of the order of $10^{-7}-10^{-5}$, the existence of non-trivial solutions to this condition implies that momenta $q$ are as large as possible, suggesting that $q_{1}\approx q_{2}\approx k/2$. We denote this type of processes as large momentum transfer and illustrate them in Fig. 3(a).
@@
Figure 4: Decay channel plot of mode $k=150$ assuming $k,p\to q_{1},q_{2}$, and setting $T=0.1\,\text{K}$ and $\kappa^{0}/(2\pi)=5\,\text{MHz}$. The darker regions correspond to participating decay channels. The colored frames highlight two relevant and distinguished types of scattering processes: with large momentum transfer (i) (light and dark blue) and small momentum transfer (ii) (brown), respectively. Numerically, for the resonant processes of type (i) (see the main text), we sum $p$ over $30\leq p\leq-1$, which includes $p^{*}$ up to $k\approx 450$. More details on the selected channels in the other cases are discussed in the main text.
@@ III.2 Resonant scattering, large momentum transfer
A non-vanishing on-shell contribution requires the existence of mode $p^{*}$ as in Eq. (33), i.e. $p^{*}$ must be larger or equal to one, resulting in the condition $k\geq\sqrt[3]{{8}/({3\xi})}$Here we used $q_{1}=q_{2}=k/2$, which gives the largest possible $p^{*}$. For the considered chain parameters, this yields $k\geq 150$. In addition, the mode with index $p^{*}$ must be occupied by at least one excitation to considerably contribute to the decay process, suggesting that temperature should be sufficiently large. However, provided that we are typically concerned with modes with small number, $p^{*}$ of order up to 10, and that typical temperatures lead to occupation of (several) tens of modes, this condition does not represent a considerable restriction.
@@
We note, that replacing the broadened density of states by a delta-function to obtain the on-shell contribution is justified if at least a few modes lie within the Lorentzian width, which is $2\kappa^{0}/v$ after linearization. This requires the inter-mode spacing given by $v$ to be smaller than the mode width $\kappa^{0}$, $\kappa^{0}>v/2$. This condition practically does not hold, so this approximation is not valid. Nevertheless, let us give for completeness a closed expression for $\delta\kappa_{k}^{(i),\text{on}}$, that is obtained from approximately replacing sums by integrals:
@@
This expression matches the result of Ref. [32], provided that periodic boundary conditions are replaced by open boundaries and notations are matched.
@@
Figure 5: Analysis of decays arising from process (i). (a) $\delta\kappa/2\pi,\delta\kappa^{(i),\text{off}}/2\pi,\delta\kappa^{(i),\text{on}}/2\pi$ as a function of mode number, for $T=0.002\,\text{K}$. The on-shell contributions strongly underestimate the total decay rate, although well-obeying the predicted power law $\sim k^{4}$ scaling with mode number. (b) The same quantities for fixed mode $k=180$, as a function of temperature. Above $T=0.005\,\text{K}$, small momentum transfer processes start to contribute to $\delta\kappa$, causing the deviation between the full decay rate and its estimate from large momentum transfer processes only. The contribution from off-resonant large momentum processes follows the expected $\sim T^{2}$ scaling, the contribution from resonant processes shows a weaker scaling although not exactly $\sim T$ (not all conditions for such scaling are fulfilled).
@@ IV.2 Enhancing resonant scattering with weak pumping
The additional peaks highlighted by blue lines in Fig. 7(a) become more prominent as intensity is increased to $\alpha=5$, however for even larger intensity $\alpha=10$ they become less visible as the distribution function strongly deviates from the Bose-Einstein distribution function everywhere. The location of blue lines is obtained independently of the solution of kinetic equation, to show resonant scattering processes. To this end, we consider all 2-into-2 scattering events $k,p\to q_{1},q_{2}$, where two modes out of $\{p,q_{1},q_{2}\}$ belong to the set of driven modes $\mathbb{K}\cup\{k^{\text{pump}}\}$. Among those combinations, we highlight mode $k$, provided the energy difference is zero within half a mode linewidth $\kappa^{0}/2$ to allow for a very small displace from resonance, $|\omega(k)+\omega(p)-\omega(q_{1})-\omega(q_{2})|\leq{\kappa^{0}}/2$. For $\omega(k)$ we use the full expression in Eq. (6). For the specific pumping configuration considered here, we get resonant in-scattering into modes k=151,152,189,220,246,270{k=151,152,189,220,246,270} where blue peaks are shown. These agree well with the deviation of $n^{0}_{k}$ from the thermal equilibrium value, except for the first two where no enhancement is noticeable. The absence of the enhancement of $n^{0}_{k}$ for smallest values of $k$ in the set of resonant modes may be attributed to the matrix element, that increases with mode numbers.
@@ V Summary and discussion
In the first part of this work, we discussed the impact of a finite linewidth and the discreteness of plasmonic modes on their lifetime at thermal equilibrium. We first connected to previous literature which assumed a continuum of zero-linewidth modes [31, 32], and discussed in which limit their results are recovered. These works considered the only possible on-shell process, which involves a large momentum difference between initial and final states and leads to an excess linewidth $\delta\kappa_{k}\sim Tk^{4}$. In our work, we showed that such process is hard to observe once modes are discrete and broadened, as a stringent set of conditions is needed. Due to broadening, deviations from such resonance are allowed, changing the expected scaling to $\delta\kappa_{k}\sim T^{2}k^{4}$. However, the resulting decay rate is predicted to be smaller than bare linewidth and therefore still challenging to observe experimentally. More importantly, broadening allows for another class of processes which would otherwise be off-resonant, with a small momentum difference between initial and final states. This features a power law $\delta\kappa_{k}\sim T^{3}k^{2}$, and can be comparable to bare linewidth at high temperature and mode number. In order to give an intuitive explanation of all scalings, we provided closed-form analytical expressions which match scalings predicted by numerical results.
@@ Acknowledgements.
We acknowledge useful discussions with Anton Bubis, Thomas Scaffidi, Alexander Mirlin, and Vladimir Manucharyan. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), and by the Austrian Science Fund (FWF) SFB F86.

### 03 Apr 2026 v2

# Plasmon decay and non-equilibrium steady states in Josephson junction chains

@@ I Introduction
Figure 1: (a) Sketch of system, consisting of a chain of JJs coupled to a transmission line. Small squares denote adjacent superconducting islands. Pink and cyan are used to distinguish adjacent islands, as in (b). Lines in different shades of magenta color illustrate the wave function of the first plasmon modes with number $k=1,2,3$. The coupling to terminals $\kappa_{\text{ex}}$ and to the internal environment $\kappa_{\text{i}}$ are also shown. (b) Detail of the circuit, with $E_{J}$ the Josephson energy, $C_{g}$ the capacitance to ground, and $\theta_{i}$ the superconducting phase of each island. (c) Comparing exact, Eq. (6), and expanded, Eq. (11) dispersion relation of the chain shows agreement for low-lying modes.
@@ II.1 From the non-interacting Hamiltonian to four-wave mixing
The first term is the charging energy and the second one is the nonlinear Josephson potential, with $E_{J}=(\hbar/2e)^{2}/L_{J}$, $L_{J}$ being the inductance of a single junction. The two variables $Q_{n}$ and $\theta_{n}$ are canonically conjugated, $\theta_{n},Q_{m}]=\mathrm{i}2e\delta_{n,m}$. Due to the low impedance of the external circuit, the phase is effectively pinned at the boundaries, and we impose $\theta_{0}=\theta_{N}=0$. Notice that such constraint is responsible for the limits 1 and N2N-2 of the second sum in Eq. (1). Regarding the charging part, $Q_{n}$ are the canonical variables conjugate to the phases, but it is sufficient to define them only for the phases that are actual dynamical degrees of freedom, namely $\theta_{1},\ldots,\theta_{N-1}$, given the others are set to zero. For this reason, the charging sum runs from $1$ to N1N-1. In the simplest case, each junction is only coupled to ground with on-site capacitance $C_{g}$ and to its nearest neighbors with capacitance $C_{J}$, such that the capacitance matrix $C_{nm}$ is tridiagonal,
@@
In the remainder of the paper, we will neglect Kerr corrections and rather focus on four-wave mixing. We restrict our analysis to number-conserving terms, as previous works have shown them to play a dominant role in the relaxation of bosons in similar contexts [48, 49]. Moreover, two-into-two processes admit on-shell configurations, i.e., channels where momentum and energy resonant conditions are satisfied even in the absence of modes’ broadening. Slight deviations from these configurations lead to only weak energy mismatch, compatible with the finite linewidth of the modes. In contrast, one-into-three scattering processes do not have on-shell channels, resulting in unfavorable, strongly off-resonant conditions. More details can be found in Appendix C.
@@ II.2 Kinetic equation and decay rates
To describe the scattering processes from quartic nonlinearity, we use the framework of a kinetic equation [50, 51, 52]. The kinetic equation provides a simplified description of the interacting modes, that neglects coherences between different modes (justified, in particular, since we are interested in steady state properties and coupling to environment destroys phase coherence) and assumes that the state of the system is specified by the occupation numbers of individual modes,
@@ II.3 Relevant parameters and cavity language
Parameter Scaling Typical value
$\mathcal{F}$ $\frac{\sqrt{E_{g}E_{J}}}{N\kappa^{0}}$ $\approx 14$
$\zeta$ $\frac{kE_{g}}{N^{2}E_{c}}$ $\sim 10^{-5}-10^{-3}$
$\bar{p}$ (w/o drive) $\frac{TN}{\sqrt{E_{g}E_{J}}}$ $\approx 30$
$g$ $\frac{1}{N^{2}}\sqrt{\frac{E_{g}}{E_{J}}}$ $\approx 1.8\cdot 10^{-8}$
Table 2: Parameters that determine the properties of a JJ chain viewed as a multi-mode cavity and their dependence on characteristics of individual $J$s, such as their charging ($E_{g}$, $E_{c}$) and Josephson ($E_{J}$) energies, as well as number of junctions NN, total coupling to the environment $\kappa^{0}$ and mode number, $k$. When appearing in the above estimates, we use temperature $T=0.1\,\text{K}$ and mode numbers in the range k=2200k=2-200.
@@
[…]
The parameters discussed before can be thought as intrinsic properties of JJ chains, as they become strongly modified only when the system is driven far away from equilibrium state. Now, we discuss two parameters that are strongly influenced by the presence of driving. The first parameter is the (effective) temperature in units of $\Delta_{1}$, that determines how many modes are occupied by at least one photon,
@@ III.1 Different processes contributing to decays
where an expansion assuming the dispersion’s nonlinearity parameter $\xi\ll 1$ (see Eq. (11)) was performed. For realistic values of parameter $\xi$, that are typically of the order of $10^{-7}-10^{-5}$, the existence of non-trivial solutions to this condition implies that momenta $q$ are as large as possible, suggesting that $q_{1}\approx q_{2}\approx k/2$. We denote this type of processes as large momentum transfer and illustrate them in Fig. 3(a).
@@
Figure 4: Decay channel plot of mode $k=150$ assuming $k,p\to q_{1},q_{2}$, and setting $T=0.1\,\text{K}$ and $\kappa^{0}/(2\pi)=5\,\text{MHz}$. The darker regions correspond to participating decay channels. The colored frames highlight two relevant and distinguished types of scattering processes: with large momentum transfer (i) (light and dark blue) and small momentum transfer (ii) (brown), respectively. Numerically, for the resonant processes of type (i) (see the main text), we sum $p$ over $30\leq p\leq-1$, which includes $p^{*}$ up to $k\approx 450$, and we restrict to $q_{1},q_{2}>0$. For the off-resonant counterpart, we consider $p\in(0,T/v]$ and $k/4\leq q_{1},q_{2}\leq 3k/4$, which correspond to a large momentum transfer window. Processes of type (ii) are instead selected through $q_{1},q_{2}>0$ and $k-\delta<q_{1},q_{2}<k+\delta$. The choice of $\delta$ and more details on the selected channels are discussed in the main text.
@@ III.2 Resonant scattering, large momentum transfer
A non-vanishing on-shell contribution requires the existence of mode $p^{*}$ as in Eq. (33), i.e. $|p^{*}|$ must be larger or equal to one, resulting in the condition $k\geq\sqrt[3]{{8}/({3\xi})}$Here we used $q_{1}=q_{2}=k/2$, which gives the largest possible $p^{*}$. For the considered chain parameters, this yields $k\geq 150$. In addition, the mode with index $p^{*}$ must be occupied by at least one excitation to considerably contribute to the decay process, suggesting that temperature should be sufficiently large. However, provided that we are typically concerned with modes $k$ in a range $\sim 1-10\,\text{GHz}$, with our chain parameters the resulting $|p^{*}|$ is of order up to $\sim 30$. Typical temperatures of order $0.1\,\text{K}$ lead to the occupation of (several) tens of modes, so this condition does not represent a considerable restriction.
@@
We note, that replacing the broadened density of states by a delta-function to obtain the on-shell contribution is justified if at least a few modes lie within the Lorentzian width, which is $2\kappa^{0}/v$ after linearization. This requires the inter-mode spacing given by $v$ to be smaller than the mode width $\kappa^{0}$, $\kappa^{0}>v/2$. This condition practically does not hold, so the approximation is not valid. Nevertheless, let us give for completeness a closed expression for $\delta\kappa_{k}^{(i),\text{on}}$, that is obtained from approximately replacing sums by integrals:
@@
This expression matches the result of Ref. [32], provided that periodic boundary conditions are replaced by open boundaries and notations are matched. A violation of $\kappa^{0}>v/2$ would affects in particular the temperature scaling, which is expected to be higher than that predicted by Eq. (37).
@@
Figure 5: Analysis of decays arising from process (i), involving a large momentum transfer between the two initial and the two final states. (a) Excess linewidth $\delta\kappa/2\pi,\delta\kappa^{(i),\text{off}}/2\pi,\delta\kappa^{(i),\text{on}}/2\pi$ as a function of mode number, for $T=0.002\,\text{K}$. The on-shell contributions $\delta\kappa^{(i),\text{on}}/2\pi$ strongly underestimate the total decay rate $\delta\kappa/2\pi$, although well-obeying the predicted power law $\sim k^{4}$ scaling with mode number. (b) The same quantities for fixed mode $k=180$, as a function of temperature. Above $T=0.005\,\text{K}$, small momentum transfer processes start to contribute to $\delta\kappa$, causing the deviation between the full decay rate and its estimate from large momentum transfer processes only. The contribution from off-resonant large momentum processes, $\delta\kappa^{(i),\text{off}}/2\pi$, follows the expected $\sim T^{2}$ scaling, while the contribution from resonant processes shows a weaker scaling although not exactly $\sim T$ (not all conditions for such scaling are fulfilled).
@@ IV.2 Enhancing resonant scattering with weak pumping
The additional peaks highlighted by blue lines in Fig. 7(a) become more prominent as intensity is increased to $\alpha=5$, however for even larger intensity $\alpha=10$ they become less visible as the distribution function strongly deviates from the Bose-Einstein distribution function everywhere. The location of blue lines is obtained independently of the solution of kinetic equation, to show resonant scattering processes. To this end, we consider all 2-into-2 scattering events $k,p\to q_{1},q_{2}$, where two modes out of $\{p,q_{1},q_{2}\}$ belong to the set of driven modes $\mathbb{K}\cup\{k^{\text{pump}}\}$. Among those combinations, we highlight mode $k$, provided the energy difference is zero within half a mode linewidth $\kappa^{0}/2$ to allow for a very small displace from resonance, $|\omega(k)+\omega(p)-\omega(q_{1})-\omega(q_{2})|\leq{\kappa^{0}}/2$. For $\omega(k)$ we use the full expression in Eq. (6). For the specific pumping configuration considered here, we get resonant in-scattering into modes k=151,152,189,220,246,270{k=151,152,189,220,246,270} where blue peaks are shown. We find that these peaks correspond to resonant scattering with large momentum transfer, where $k^{\text{pump}}$ scatters off the negative component of a pumped mode in $\mathbb{K}$, $p^{*}=-1,-2,-3,-4,-5$ (see the schematic of Fig. 3(a)). These agree well with the deviation of $n^{0}_{k}$ from the thermal equilibrium value, except for the first two where no enhancement is noticeable. The absence of the enhancement of $n^{0}_{k}$ for smallest values of $k$ in the set of resonant modes may be attributed to the matrix element, that increases with mode numbers.
@@ V Summary and discussion
In the first part of this work, we discussed the impact of a finite linewidth and the discreteness of plasmon modes on their lifetime at thermal equilibrium. We first connected to previous literature which assumed a continuum of zero-linewidth modes [31, 32], and discussed in which limit their results are recovered. These works considered the only possible on-shell process, which involves a large momentum difference between initial and final states and leads to an excess linewidth $\delta\kappa_{k}\sim Tk^{4}$. In our work, we showed that such process is hard to observe once modes are discrete and broadened, as a stringent set of conditions is needed. Due to broadening, deviations from such resonance are allowed, changing the expected scaling to $\delta\kappa_{k}\sim T^{2}k^{4}$. However, the resulting decay rate is predicted to be smaller than bare linewidth and therefore still challenging to observe experimentally. More importantly, broadening allows for another class of processes which would otherwise be off-resonant, with a small momentum difference between initial and final states. This features a power law $\delta\kappa_{k}\sim T^{3}k^{2}$, and can be comparable to bare linewidth at high temperature and mode number. In order to give an intuitive explanation of all scalings, we provided closed-form analytical expressions which match scalings predicted by numerical results.
@@ Acknowledgements.
We acknowledge useful discussions with Anton Bubis, Vladimir Manucharyan, Alexander Mirlin and Thomas Scaffidi. We are grateful to Eli Levenson-Falk and Alexander Mirlin for their feedback on the manuscript. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), and by the Austrian Science Fund (FWF) SFB F86.